The dimeric compound [(n5-C5Me5)Cr(CO)2]2 (A) follows rule 18e and in the IR spectrum shows absorptions at 1870 cm‒1. Treatment of the compound with excess CO at 1200-450 psi and 175oC for 10 hours gives another dimer compound of chromium (B) which also follows rule 18e and shows infrared absorptions at 1876 cm‒1. Photolysis of compound A by UV radiation leads to dimer compound C together with the release of CO gas. The IR spectrum of compound C shows only one band at 1788 cm‒1 while also showing the same metal-metal bond order as compound A. Propose suitable molecular structures for compounds A, B and C by clearly indicating the Cr την bond order Cr and how to assemble the CO.


1
Expert's answer
2021-02-18T07:28:26-0500

The anodic electrochemical oxidations of ReCp(CO)3 (1, Cp = eta(5)-C5H5), Re(eta(5)-C5H4NH2)(CO)3 (2), and ReCp*(CO)3 (3, Cp* = eta(5)-C5Me5), have been studied in CH2Cl2 containing [NBu4][TFAB] (TFAB = [B(C6F5)4]-) as supporting electrolyte. One-electron oxidations were observed with E(1/2) = 1.16, 0.79, and 0.91 V vs ferrocene for 1-3, respectively. In each case, rapid dimerization of the radical cation gave the dimer dication, [Re2Cp(gamma)2(CO)6]2+ (where Cp(gamma) represents a generic cyclopentadienyl ligand), which may be itself reduced cathodically back to the original 18-electron neutral complex ReCp(gamma)(CO)3. DFT calculations show that the SOMO of 1+ is highly Re-based and hybridized to point away from the metal, thereby facilitating the dimerization process and other reactions of the Re(II) center. The dimers, isolated in all three cases, have long metal-metal bonds that are unsupported by bridging ligands, the bond lengths being calculated as 3.229 A for [Re2Cp2(CO)6]2+ (1(2)2+) and measured as 3.1097 A for [Re2(C5H4NH2)2(CO)6]2+ (2(2)2+) by X-ray crystallography on [Re2(C5H4NH2)2(CO)6][TFAB]2. The monomer/dimer equilibrium constants are between K(dim) = 10(5) M(-1) and 10(7) M(-1) for these systems, so that partial dissociation of the dimers gives a modest amount of the corresponding monomer that is free to undergo radical cation reactions. The radical 1+ slowly abstracts a chlorine atom from dichloromethane to give the 18-electron complex [ReCp(CO)3Cl]+ as a side product. The radical cation 1+ acts as a powerful one-electron oxidant capable of effectively driving outer-sphere electron-transfer reactions with reagents having potentials of up to 0.9 V vs ferrocene.